# PALM: Pushing Adaptive Learning Rate Mechanisms for Continual Test-Time Adaptation

### v3

@@ 3 Preliminary
Core motivations: We provide a detailed view of surgical fine-tuning (Lee et al. 2022), AutoLR (Ro and Choi 2021), and LAW (Park et al. 2024). It is critical to understand the design principles and reasons for having an adaptive LR scheme for CTTA, which motivates this work.
To preserve the information during the source pre-training phase while fine-tuning on a target task of a different distribution like an image corruption, surgical fine-tuning empirically shows that heuristically fine-tuning only the first convolutional block results in better performance over full fine-tuning. Even though it is a desirable outcome for efficient fine-tuning, in this work, we close the gap by automatically selecting the layers that would adapt well to the target task.
@@ 4 Proposed Methodology
Layer selection by quantifying prediction uncertainty: CTTA maximizes model performance in the current domain and so it is vital to assess prediction certainty amidst rapid domain shifts. In the literature on continual learning and TTA, several works utilize FIM (Kirkpatrick et al. 2017; Schwarz et al. 2018; Park et al. 2024). However, the unavailability of true labels, with unknown distribution, pushes us to capture the model prediction uncertainty and adapt selected layers, automatically chosen, to a batch of test data, unlike surgical fine-tuning (Lee et al. 2022).
@@
Parameter sensitivity as a domain shift indicator: New studies have found that model parameters demonstrate levels of criticality (Chatterji, Neyshabur, and Sedghi 2019; Zhang, Bengio, and Singer 2022) across different positions. Moreover, not all parameters adapt uniformly, and removing some can lower test generalization errors (Mozer and Smolensky 1988; Rasmussen and Ghahramani 2000). Inspired by this observation, we assert that parameters contribute differently during CTTA. From the previous stage, we automatically select the layers based on the prediction uncertainty. For parameters of these selected layers only, we compute their sensitivity i.e., an approximation of an error when it is removed. Motivated by (LeCun, Denker, and Solla 1989; Molchanov et al. 2016, 2019), we compute the sensitivity of a parameter $\theta_{j,n}^{t}$ based on the loss (Eq. 1) as follows,
@@ 5 Experiments and Results
BN-1 (Schneider et al. 2020) 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
TENT Continual (Wang et al. 2020) 24.8 20.6 28.6 14.4 31.1 16.5 14.1 19.1 18.6 18.6 12.2 20.3 25.7 20.8 24.9 20.7
CoTTA (Wang et al. 2022)
24.3 21.3 26.6 11.6 27.6 12.2 10.3 14.8 14.1 12.4 7.5 10.6 18.3 13.4 17.3 16.2
EcoTTA (Song et al. 2023)
23.8 18.7 25.7 11.5 29.8 13.3 11.3 15.3 15.0 13.0 7.9 11.3 20.2 15.1 20.5 16.8
AdaContrast (Chen et al. 2022)
29.1 22.5 30.0 14.0 32.7 14.1 12.0 16.6 14.9 14.4 8.1 10.0 21.9 17.7 20.0 18.5
Surgical
fine-tuning (Lee et al. 2022)
27.7 24.1 33.8 12.2 33.1 13.1 10.9 15.9 15.6 13.8 7.8 10.8 21.2 16.6 22.9 18.6
DSS (Wang et al. 2024)
24.1 21.3 25.4 11.7 26.9 12.2 10.5 14.5 14.1 12.5 7.8 10.8 18.0 13.1 17.3 16.0
LAW (Park et al. 2024)
24.7 18.9 25.5 12.9 26.7 15.0 11.8 15.1 14.7 15.9 10.1 13.8 19.4 14.7 18.3 17.2
Ours
25.8 18.1 22.7 12.3 25.3 13.1 10.7 13.5 13.1 12.2 8.5 11.8 17.9 12.0 15.4 15.5
@@
Source
73.0 68 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
BN-1 (Schneider et al. 2020)
42.1 40.7 42.7 27.6 41.9 29.7 27.9 34.9 35.0 41.5 26.5 30.3 35.7 32.9 41.2 35.4
TENT Continual (Wang et al. 2020)
37.2 35.8 41.7 37.9 51.2 48.3 48.5 58.4 63.7 71.1 70.4 82.3 88.0 88.5 90.4 60.9
CoTTA (Wang et al. 2022)
40.1 37.7 39.7 26.9 38.0 27.9 26.4 32.8 31.8 40.3 24.7 26.9 32.5 28.3 33.5 32.5
AdaContrast (Chen et al. 2022)
42.3 36.8 38.6 27.7 40.1 29.1 27.5 32.9 30.7 38.2 25.9 28.3 33.9 33.3 36.2 33.4
Surgical
fine-tuning (Lee et al. 2022)
40.7 36.3 37.8 25.5 39.0 27.6 25.2 32.2 30.4 37.3 23.7 26.8 32.3 27.6 37.2 32.0
DSS(Wang et al. 2024)
39.7 36.0 37.2 26.3 35.6 27.5 25.1 31.4 30.0 37.8 24.2 26.0 30.0 26.3 31.1 30.9
LAW (Park et al. 2024)
41.0 36.7 38.3 25.6 37.0 27.8 25.2 30.7 30.0 37.3 24.4 27.6 31.2 27.8 34.9 31.7
Ours
37.3 32.5 34.9 26.2 35.3 27.5 24.6 28.8 29.1 34.1 23.5 26.9 31.2 26.6 34.1 30.1
@@
CTTA implementation details: To maintain fairness regarding the source models, we employ WideResNet-28 (Zagoruyko and Komodakis 2016) for CIFAR-10C, ResNeXt-29 (Xie et al. 2017) for CIFAR-100C, and ResNet-50 (He et al. 2016) for ImageNet-C, all available on RobustBench (Croce et al. 2020). We optimize using an Adam optimizer, setting base learning rates ($\kappa$) to 5e-4 for CIFAR-10C and CIFAR-100C, and 5e-5 for ImageNet-C. For balancing parameter sensitivities, we set $\alpha$ to 0.5, 0.9, and 0.5 respectively, with temperature coefficients T set to 50, 100, and 1000 respectively. We set $\eta$ to 1, 0.5, and 0.3, and $\lambda$ to 0.01 throughout. Batch sizes are set to 200, 200, and 64 for each dataset, following CoTTA, and results are only reported for the highest severity level of 5 for each task.
@@
Baselines: We compare against several CTTA baselines, including BN Stats Adapt (BN-1) (Schneider et al. 2020), TENT Continual (Wang et al. 2020), CoTTA (Wang et al. 2022), EcoTTA (Song et al. 2023), AdaContrast (Chen et al. 2022), DSS (Wang et al. 2024), and LAW (Park et al. 2024). We also contrast our work against surgical fine-tuning (Lee et al. 2022). For EcoTTA, we report the results on CIFAR-10C and ImageNet-C only since they do not use ResNeXt-29 (Xie et al. 2017) for CIFAR-100C. Following the findings of (Lee et al. 2022), we continually fine-tune only the first convolutional block of the respective models, denoted as “Surgical fine-tuning (continual)”, using an SGD optimizer with fixed learning rates of 1e-3 for the CIFAR datasets and 5e-4 for ImageNet-C. For the results of LAW, we reproduce their results strictly following their official implementation  [^1] https://github.com/junia3/LayerwiseTTA/tree/main with parameters as mentioned in their paper.
@@
Source
97.8 97.1 98.2 81.7 89.8 85.2 78.0 83.5 77.1 75.9 41.3 94.5 82.5 79.3 68.6 82.0
BN-1 (Schneider et al. 2020)
85.0 83.7 85.0 84.7 84.3 73.7 61.2 66.0 68.2 52.1 34.9 82.7 55.9 51.3 59.8 68.6
TENT Continual (Wang et al. 2020)
81.6 74.6 72.7 77.6 73.8 65.5 55.3 61.6 63.0 51.7 38.2 72.1 50.8 47.4 53.3 62.6
CoTTA (Wang et al. 2022)
84.7 82.1 80.6 81.3 79.0 68.6 57.5 60.3 60.5 48.3 36.6 66.1 47.2 41.2 46.0 62.7
EcoTTA (Song et al. 2023)
- - - - - - - - - - - - - - - 63.4
AdaContrast (Chen et al. 2022)
82.9 80.9 78.4 81.4 78.7 72.9 64.0 63.5 64.5 53.5 38.4 66.7 54.6 49.4 53.0 65.5
Surgical
fine-tuning (Lee et al. 2022)
83.4 79.5 77.2 86.1 82.4 72.2 61.1 63.1 64.7 51.2 35.9 72.9 56.2 46.7 50.2 65.5
DSS (Wang et al. 2024)
84.6 80.4 78.7 83.9 79.8 74.9 62.9 62.8 62.9 49.7 37.4 71.0 49.5 42.9 48.2 64.6
LAW (Park et al. 2024)
81.8 75.0 72.1 77.4 73.2 63.9 54.6 57.9 61.0 50.1 36.3 68.6 49.0 46.2 49.1 61.0
Ours
81.1 73.3 70.8 77.0 71.8 62.3 53.9 56.7 60.8 50.4 36.2 65.9 48.0 45.2 48.0 60.1
@@
Results on CIFAR-10C and CIFAR-100C: We report the CTTA results in Tables 1 and 2. Our method performs the best across all training setups and baselines for both datasets. We achieve large improvements of 4.9% and 5.2% for CIFAR-10C and 5.3% and 30.8% for CIFAR-100C, over BN-1 and TENT Continual respectively. Concerning recent works of LAW and DSS, our method improves by 1.7% and 0.5% on CIFAR-10C and 1.6% and 0.8% on CIFAR-100C, respectively. TENT Continual excels initially but accumulates errors over time, increasing mean error. In CoTTA, the method’s reliance on full model updates renders it computationally expensive and unsuitable for deployment. An interesting observation is that with a drastic reduction in the number of trainable parameters, as in “Surgical fine-tuning (continual)”, the mean errors are better than most CTTA benchmarks, if not the best. LAW uses the uncertain pseudo labels to calculate the domain shift approximated via the FIM. The minor error that is introduced for each batch begins to propagate across all the domains. In addition, the imbalance caused by the “domain-level” FIM and the current batch’s FIM, results in a higher error rate for the initial tasks. In our method, the computation of the model prediction uncertainty to automatically select layers is useful (Eq. 1). This affirms the findings in (Lee et al. 2022) that concentrated adaptation, for image corruptions, in the initial layers suffices.
Results on Imagenet-C: In Table 3, we report the CTTA results on the ImageNet-C dataset. As shown, our method outperforms all the previous CTTA baselines and “Surgical fine-tuning (continual)”. We observe an improvement of 0.9% and 4.5% over LAW and DSS respectively. In addition, we achieve 2.6%, 5.4%, and 2.5% improvements over CoTTA, AdaContrast, and TENT Continual, respectively.
@@
Table 4: Mean errors (in %) on CIFAR-10C, CIFAR-100C, and ImageNet-C - We report the GTTA mean errors averaged across all the severity levels (1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5).
@@
Computational efficiency and Catastrophic forgetting (CF): In our PALM, we first select a subset of layers from the gradient space (Eq. 2) and adapt them while strictly freezing the unselected layers. In Table 5, we report the average % of adapted model parameters per task with respect to the source model, for our method and CoTTA, across the benchmark datasets. Our method is computationally inexpensive and on an average updates 4.34%, 6.65%, and 17.21% of their respective source parameters. This showcases significant computational advantages over CoTTA, which updates 100% of the parameters. With a single NVIDIA A5000 GPU, PALM incurs a slightly higher adaptation time/batch, due to the two-stage proposed method. For example, on CIFAR-10, PALM runs for 0.8s per batch, whereas Surgical Fine-tuning (Surgical) and LAW take 0.45 s and 0.73 s respectively. Our method alleviates CF by only adapting a small subset of model parameters, ensuring that much of the prior source knowledge is preserved. Quantitatively, the source models achieved zero-shot mean accuracies of 94.03%, 76.06%, and 75.19% on respective datasets. After adaptation, PALM retained source knowledge better, with mean accuracies of 92.29%, 77.46%, and 70.87%, compared to LAW’s 91.67%, 77.69%, and 71.57%. Despite accuracy decline with more tasks, especially after adaptation on ImageNet-C, PALM matches or outperforms LAW.
Dataset Method
The average % of
adapted parameters/task
(source parameters)
CIFAR-10C Ours 4.34 (of 36.4M)
CoTTA 100 (of 36.4M)
CIFAR-100C Ours 6.65 (of 6.8M)
CoTTA 100 (of 6.8M)
ImageNet-C Ours 17.21 (of 23.5M)
CoTTA 100 (of 23.5M)
Table 5: Average % of adapted parameters/task.
Adaptive LR optimizers vs Ours: Here, we study the effect of directly using adaptive LR optimizers like LARS (You, Gitman, and Ginsburg 2017), Prodigy (Mishchenko and Defazio 2023), and Adafactor (Shazeer and Stern 2018), for CTTA. We report the results in Table 6. The mentioned optimizers do not adjust layer-wise LRs based on parameter sensitivity, failing to capture rapid domain shifts, resulting in poorer performance compared to our method.
@@
Figure 2: Illustrations of LR importance $i_{j,n}^{t}$ across different convolutional blocks/stages/layers during CTTA. [Top row] - Variations for the domain “glass_blur” across datasets. [Bottom row] - Variations for the domain “snow”.
@@
Refer to caption Refer to caption
a) Ablation on $\alpha$. b) Ablation on temperature T.
@@
Table 7: Layer importance using pseudo-labels vs ours.
@@
Contribution of smoothing factor $\alpha$: A key component of PALM is computing the moving average of sensitivities across data batches, smoothed by $\alpha$ (Eq. 5). Figure 3(a) illustrates the variations of the mean errors, for the benchmark datasets, for varying $\alpha$. We observe that the “domain-level” sensitivity across past batches of data is indeed useful. For $\alpha$ = 0 i.e., when $S_{j,n}^{t}$ of a selected parameter at timestep t is not considered, the respective datasets’ error rate drastically spikes. For CIFAR-10C, we achieve the best performance when $\alpha$ = 0.5 with an improvement of 1.7% over LAW. For CIFAR-100C and ImageNet-C, we see improvements of 1.6% and 0.9% over LAW, at $\alpha$ = 0.9 and 0.5 respectively. With higher values of $\alpha$, the results do not vary too much. Hence, attaining a balance between sensitivities from past domains ($\hat{S}_{j,n}^{t}$) and current sensitivity ($S_{j,n}^{t}$) is needed.
@@
Effectiveness of temperature T: In our method, we compute the KL divergence between a uniform distribution u and a smoothed classifier’s output logits (Eq. 1), denoted by $\hat{h}(x_{k})$. To study its influence, we illustrate the results in Figure 3(b). Ideally, larger values of T push the distributions of the predictions closer to u (Hinton, Vinyals, and Dean 2015) by spreading the probability among inputs. In CTTA, since we deal with domain shifts with no labels, we argue that higher values of T are useful to capture the uncertainty. So, the KL divergence reduces, indicating more uncertainty in the model and hence, more adaptation. For a systematic temperature selection, we aim to adjust T based on the number of adapted parameters. In particular, if the number of adapted parameters increases across batches, T will be increased. For a difficult dataset like ImageNet-C, larger values of T are useful with the error reducing as T increases. For the CIFAR-C datasets, the errors begin to rise for T $\geq$ 100 which could be due to large spreads of the distributions.
@@
Robustness to regularization $\lambda$: Here, we understand the effect of the parameter $\lambda$ in Eq. 11. We set $\lambda$in {0, 0.01, 0.05, 0.1, 0.5, 1.0}. Table 8 reports the average classification errors on the benchmark datasets. For $\lambda$ = 0.01, we achieve the best results. In addition, we observe that our method is very robust to $\lambda$ i.e., with subtle variations to differing $\lambda$.
6 Conclusion

### v4

@@ Introduction
Core motivations: We provide a detailed view of surgical fine-tuning (SFT) (Lee et al. 2022), AutoLR (Ro and Choi 2021), and LAW (Park et al. 2024). It is critical to understand the design principles and reasons for having an adaptive LR scheme for CTTA, which motivates this work.
To preserve the information during the source pre-training phase while fine-tuning on a target task of a different distribution like an image corruption, SFT empirically shows that heuristically fine-tuning only the first convolutional block results in better performance over full fine-tuning. Even though it is a desirable outcome for efficient fine-tuning, in this work, we close the gap by automatically selecting the layers that would adapt well to the target task.
@@
Layer selection by quantifying prediction uncertainty: CTTA maximizes model performance in the current domain and so it is vital to assess prediction certainty amidst rapid domain shifts. In the literature on continual learning and TTA, several works utilize FIM (Kirkpatrick et al. 2017; Schwarz et al. 2018; Park et al. 2024). However, the unavailability of true labels, with unknown distribution, pushes us to capture the model prediction uncertainty and adapt selected layers, automatically chosen, to a batch of test data, unlike SFT (Lee et al. 2022).
@@
Parameter sensitivity as a domain shift indicator: New studies have found that model parameters demonstrate levels of criticality (Zhang, Bengio, and Singer 2022) across different positions. Moreover, not all parameters adapt uniformly, and removing some can lower test generalization errors (Mozer and Smolensky 1988; Rasmussen and Ghahramani 2000). Inspired by this observation, we assert that parameters contribute differently during CTTA. From the previous stage, we automatically select the layers based on the prediction uncertainty. For parameters of these selected layers only, we compute their sensitivity i.e., an approximation of an error when it is removed. Motivated by (Molchanov et al. 2016, 2019), we compute the sensitivity of a parameter $\theta_{j,n}^{t}$ based on the loss (Eq. 1) as follows,
@@
BN-1 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
Tent Cont. 24.8 20.6 28.6 14.4 31.1 16.5 14.1 19.1 18.6 18.6 12.2 20.3 25.7 20.8 24.9 20.7
CoTTA 24.3 21.3 26.6 11.6 27.6 12.2 10.3 14.8 14.1 12.4 7.5 10.6 18.3 13.4 17.3 16.2
EcoTTA 23.8 18.7 25.7 11.5 29.8 13.3 11.3 15.3 15.0 13.0 7.9 11.3 20.2 15.1 20.5 16.8
AdaContrast 29.1 22.5 30.0 14.0 32.7 14.1 12.0 16.6 14.9 14.4 8.1 10.0 21.9 17.7 20.0 18.5
SFT (cont.) 27.7 24.1 33.8 12.2 33.1 13.1 10.9 15.9 15.6 13.8 7.8 10.8 21.2 16.6 22.9 18.6
DSS 24.1 21.3 25.4 11.7 26.9 12.2 10.5 14.5 14.1 12.5 7.8 10.8 18.0 13.1 17.3 16.0
LAW 24.7 18.9 25.5 12.9 26.7 15.0 11.8 15.1 14.7 15.9 10.1 13.8 19.4 14.7 18.3 17.2
Ours 25.8 18.1 22.7 12.3 25.3 13.1 10.7 13.5 13.1 12.2 8.5 11.8 17.9 12.0 15.4 15.5
@@
Source 73.0 68 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
BN-1 42.1 40.7 42.7 27.6 41.9 29.7 27.9 34.9 35.0 41.5 26.5 30.3 35.7 32.9 41.2 35.4
Tent Cont. 37.2 35.8 41.7 37.9 51.2 48.3 48.5 58.4 63.7 71.1 70.4 82.3 88.0 88.5 90.4 60.9
CoTTA 40.1 37.7 39.7 26.9 38.0 27.9 26.4 32.8 31.8 40.3 24.7 26.9 32.5 28.3 33.5 32.5
AdaContrast 42.3 36.8 38.6 27.7 40.1 29.1 27.5 32.9 30.7 38.2 25.9 28.3 33.9 33.3 36.2 33.4
SFT (cont.) 40.7 36.3 37.8 25.5 39.0 27.6 25.2 32.2 30.4 37.3 23.7 26.8 32.3 27.6 37.2 32.0
DSS 39.7 36.0 37.2 26.3 35.6 27.5 25.1 31.4 30.0 37.8 24.2 26.0 30.0 26.3 31.1 30.9
LAW 41.0 36.7 38.3 25.6 37.0 27.8 25.2 30.7 30.0 37.3 24.4 27.6 31.2 27.8 34.9 31.7
Ours 37.3 32.5 34.9 26.2 35.3 27.5 24.6 28.8 29.1 34.1 23.5 26.9 31.2 26.6 34.1 30.1
@@
CTTA implementation details: To maintain fairness regarding the source models, we employ WideResNet-28 (36.4M params) (Zagoruyko and Komodakis 2016) for CIFAR-10C, ResNeXt-29 (6.8M params) (Xie et al. 2017) for CIFAR-100C, and ResNet-50 (23.5M params) (He et al. 2016) for ImageNet-C, all available on RobustBench (Croce et al. 2020). We optimize using an Adam optimizer, setting base learning rates ($\kappa$) to 5e-4 for CIFAR-10C and CIFAR-100C, and 5e-5 for ImageNet-C. For balancing parameter sensitivities, we set $\alpha$ to 0.5, 0.9, and 0.5 respectively, with temperature coefficients T set to 50, 100, and 1000 respectively. We set $\eta$ to 1, 0.5, and 0.3, and $\lambda$ to 0.01 throughout. Batch sizes are set to 200, 200, and 64 for each dataset, following CoTTA, and results are only reported for the highest severity level of 5 for each task.
@@
Baselines: We compare against several CTTA baselines, including BN Stats Adapt (BN-1) (Schneider et al. 2020), Tent Cont. (Wang et al. 2020), CoTTA (Wang et al. 2022), EcoTTA (Song et al. 2023), AdaContrast (Chen et al. 2022), DSS (Wang et al. 2024), and LAW (Park et al. 2024). We also contrast our work against surgical fine-tuning (Lee et al. 2022). For EcoTTA, we report the results on CIFAR-10C and ImageNet-C only since they do not use ResNeXt-29 (Xie et al. 2017) for CIFAR-100C. Following the findings of (Lee et al. 2022), we continually fine-tune only the first convolutional block of the respective models, denoted as “SFT (cont.)”, using an SGD optimizer with fixed learning rates of 1e-3 for the CIFAR datasets and 5e-4 for ImageNet-C. For the results of LAW, we reproduce their results strictly following their official implementation  [^1] https://github.com/junia3/LayerwiseTTA/tree/main with parameters as mentioned in their paper.
@@
Source 97.8 97.1 98.2 81.7 89.8 85.2 78.0 83.5 77.1 75.9 41.3 94.5 82.5 79.3 68.6 82.0
BN-1 85.0 83.7 85.0 84.7 84.3 73.7 61.2 66.0 68.2 52.1 34.9 82.7 55.9 51.3 59.8 68.6
Tent Cont. 81.6 74.6 72.7 77.6 73.8 65.5 55.3 61.6 63.0 51.7 38.2 72.1 50.8 47.4 53.3 62.6
CoTTA 84.7 82.1 80.6 81.3 79.0 68.6 57.5 60.3 60.5 48.3 36.6 66.1 47.2 41.2 46.0 62.7
EcoTTA - - - - - - - - - - - - - - - 63.4
AdaContrast 82.9 80.9 78.4 81.4 78.7 72.9 64.0 63.5 64.5 53.5 38.4 66.7 54.6 49.4 53.0 65.5
SFT (cont.) 83.4 79.5 77.2 86.1 82.4 72.2 61.1 63.1 64.7 51.2 35.9 72.9 56.2 46.7 50.2 65.5
DSS 84.6 80.4 78.7 83.9 79.8 74.9 62.9 62.8 62.9 49.7 37.4 71.0 49.5 42.9 48.2 64.6
LAW 81.8 75.0 72.1 77.4 73.2 63.9 54.6 57.9 61.0 50.1 36.3 68.6 49.0 46.2 49.1 61.0
Ours 81.1 73.3 70.8 77.0 71.8 62.3 53.9 56.7 60.8 50.4 36.2 65.9 48.0 45.2 48.0 60.1
@@
Results on CIFAR-10C and CIFAR-100C: We report the CTTA results in Tables 1 and 2. Our method performs the best across all training setups and baselines for both datasets. We achieve large improvements of 4.9% and 5.2% for CIFAR-10C and 5.3% and 30.8% for CIFAR-100C, over BN-1 and Tent Cont. respectively. Concerning recent works of LAW and DSS, our method improves by 1.7% and 0.5% on CIFAR-10C and 1.6% and 0.8% on CIFAR-100C, respectively. Tent Cont. excels initially but accumulates errors over time, increasing mean error. In CoTTA, the method’s reliance on full model updates renders it computationally expensive and unsuitable for deployment. An interesting observation is that with a drastic reduction in the number of trainable parameters, as in SFT (cont.), the mean errors are better than most CTTA benchmarks, if not the best. LAW uses the uncertain pseudo labels to calculate the domain shift approximated via the FIM. The minor error that is introduced for each batch begins to propagate across all the domains. In addition, the imbalance caused by the “domain-level” FIM and the current batch’s FIM, results in a higher error rate for the initial tasks. In our method, the computation of the model prediction uncertainty to automatically select layers is useful (Eq. 1). This confirms the findings in SFT (Lee et al. 2022) that concentrated adaptation, for image corruptions, in the initial layers suffices.
Results on Imagenet-C: In Table 3, we report the CTTA results on the ImageNet-C dataset. As shown, our method outperforms all the previous CTTA baselines and SFT (cont.). We observe an improvement of 0.9% and 4.5% over LAW and DSS respectively. In addition, we achieve 2.6%, 5.4%, and 2.5% improvements over CoTTA, AdaContrast, and Tent Cont., respectively.
@@
Table 4: Mean errors (in %) on CIFAR-10C, CIFAR-100C, and ImageNet-C - GTTA mean errors averaged across all severity levels (1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5).
@@
Computational efficiency and Catastrophic forgetting (CF): In our PALM, we first select a subset of layers from the gradient space (Eq. 2) and adapt them while strictly freezing the unselected layers. In Table 6, we report the average % of adapted model parameters per task with respect to the source model, for our method and CoTTA, across the benchmark datasets. Our method is computationally inexpensive and on an average updates 4.34%, 6.65%, and 17.21% of their respective source parameters. This showcases significant computational advantages over CoTTA, which updates 100% of the parameters. With a single NVIDIA A5000 GPU, PALM incurs a slightly higher adaptation time/batch, due to the two-stage proposed method. For example, on CIFAR-10, PALM runs for 0.8s per batch, whereas SFT (cont.) and LAW take 0.45 s and 0.73 s respectively. Our method alleviates CF by only adapting a small subset of model parameters, ensuring that much of the prior source knowledge is preserved. Quantitatively, the source models achieved zero-shot mean accuracies of 94.03%, 76.06%, and 75.19% on respective datasets. After adaptation, PALM retained source knowledge better, with mean accuracies of 92.29%, 77.46%, and 70.87%, compared to LAW’s 91.67%, 77.69%, and 71.57%. Despite the accuracy decline with more tasks, especially after adaptation on ImageNet-C, PALM matches or outperforms LAW.
      Dataset Ours CoTTA
CIFAR-10C 4.34 100
CIFAR-100C 6.65 100
ImageNet-C 17.21 100
Table 5: The average % of adapted source parameters/task.
@@
Adaptive LR optimizers vs Ours: Here, we study the effect of directly using adaptive LR optimizers like LARS (You, Gitman, and Ginsburg 2017), Prodigy (Mishchenko and Defazio 2023), and Adafactor (Shazeer and Stern 2018), for CTTA. We report the results in Table 6. The mentioned optimizers do not adjust layer-wise LRs based on parameter sensitivity, failing to capture rapid domain shifts and resulting in poorer performance compared to our method.
@@
Figure 2: Illustrations of LR importance $i_{j,n}^{t}$ across different convolutional blocks(B)/stages(S)/layers(L) during CTTA. [Top row] - Variations for the domain “glass_blur” across datasets. [Bottom row] - Variations for the domain “snow”.
@@
Table 7: Prediction uncertainty using pseudo-labels vs. ours.
@@
Refer to caption Refer to caption
a) Ablation on temperature T b) Ablation on $\alpha$
@@
Effectiveness of temperature T: In our method, we compute the KL divergence between the uniform distribution u and a smoothed classifier’s output logits (Eq. 1), denoted by $\hat{h}(x_{k})$. We illustrate the results in Figure 3(a) to study its influence. Ideally, larger values of T push the distributions of the predictions closer to u (Hinton, Vinyals, and Dean 2015) by spreading the probability among inputs. In CTTA, since we deal with domain shifts with no labels, we argue that higher values of T are useful to capture the uncertainty. So, the KL divergence reduces, indicating more uncertainty in the model and hence, more adaptation. For a systematic temperature selection, we aim to adjust T based on the number of adapted parameters. In particular, if the number of adapted parameters increases across batches, T will be increased. For a difficult dataset like ImageNet-C, larger values of T are useful with the error reducing as T increases. For the CIFAR-C datasets, the errors begin to rise for T $\geq$ 100 which could be due to large spreads of the distributions.
Contribution of smoothing factor $\alpha$: A key component of PALM is computing the moving average of sensitivities across data batches, smoothed by $\alpha$ (Eq. 5). Figure 3(b) illustrates the variations of the mean errors, for the benchmark datasets, for varying $\alpha$. We observe that the “domain-level” sensitivity across past batches of data is indeed useful. For $\alpha$ = 0 i.e., when $S_{j,n}^{t}$ of a selected parameter at timestep t is not considered, the respective datasets’ error rate drastically spikes. For CIFAR-10C, we achieve the best performance when $\alpha$ = 0.5 with an improvement of 1.7% over LAW. For CIFAR-100C and ImageNet-C, we see improvements of 1.6% and 0.9% over LAW, at $\alpha$ = 0.9 and 0.5 respectively. With higher values of $\alpha$, the results do not vary too much. Hence, attaining a balance between sensitivities from past domains ($\hat{S}_{j,n}^{t}$) and current sensitivity ($S_{j,n}^{t}$) is needed.
@@
Robustness to $\lambda$: In Eq. 11, we set $\lambda$in {0, 0.01, 0.05, 0.1, 0.5, 1.0}. Table 8 reports the average classification errors on the benchmark datasets. For $\lambda$ = 0.01, we achieve the best results. In addition, we observe that our method is very robust to $\lambda$ i.e., with subtle variations to differing $\lambda$.
Conclusion